Project

General

Profile

10/28/2013 Minutes

Attending: Michael, Nick, Mat, Dom, Evan, Gavin, Denis

Operations

11498 and up are good runs.

Production

  • Disk space issues. (5 TB left) (10 TB of backup to remove)

Watchdog

  • Dom reports the far detector watchdog web thinks that the processing is 6 hrs behind up to CAF
  • This webpage needs to be coordinated with Micheal
  • All shifters documented run info well except the weekend night shifter. (Mat will bring it up with Steve.)

Issues

  • The production fix is working.
  • (ADC distribution were made.)
  • Retuning slicer (Micheal)
  • NDOS timing plots missing following run name change (The suggestion is to not fix this until the final scheme is decided.)

Working time

  • Scintillator level
    Rick wants to see voids.
    Cell hit efficiency as a function of W.
    Track hits
  • Jim wants ADC versus W
  • Jim is going to tell us other items.
  • Interest in understudying the multiplexing. (Mat will get the design from the expert and look at the data.